Moles of original Ag+?

Weight of AgNO3=0.505g

Molar mass of AgNO3= 169.87 g/mol

Moles of original AgNO3= 86.10 mol
Answer #1

Moles = mass / molar mass of 1 mole AgNO3

= 0.505 / 169.87

= 0.0029 moles

Number of moles of 86.10 moles of AgNO3 = 86.1*0.003 = 0.258 moles

Now, for Ag+

Moles of Ag+ / MW of Ag = Moles of AgNO3 / MW of AgNO3

Moles of Ag = 86.1 * 107.87 / 169.87

= 54.67 moles